### v1

# Ranks of matrix factorizations and sheaf cohomology

Abstract.
Buchweitz-Greuel-Schreyer conjectured in 1987 a lower bound on the ranks of matrix factorizations over certain local hypersurface rings. We study a graded version of this conjecture, and we show that it implies a novel conjecture concerning the cohomology of sheaves over non-Fano projective hypersurfaces.
2020 Mathematics Subject Classification:
13D02, 14F08
Brown was partially supported by NSF grant DMS-2302373 and Walker by NSF grant DMS-2200732.
1. Introduction
Let $Q$ be a commutative ring and $f\in Q$. A matrix factorization of $f$ is a tuple $(F^{0},F^{1},s^{0},s^{1})$, where $F^{0}$, $F^{1}$ are finitely generated free $Q$-modules, and $s^{0}\colon F^{0}\to F^{1}$, $s^{1}\colon F^{1}\to F^{0}$ are $Q$-linear maps such that $s^{1}s^{0}=f\cdot\operatorname{id}_{F^{0}}$ and $s^{0}s^{1}=f\cdot\operatorname{id}_{F^{1}}$. A trivial matrix factorization is a direct sum of copies of $(Q,Q,1,f)$ and $(Q,Q,f,1)$. The following is a weak version of a conjecture of Buchweitz-Greuel-Schreyer:
Conjecture 1.1 (Weakening of [BGS87] Conjecture A).
Let $Q$ be a regular local ring of dimension $n+1$ with algebraically closed residue field, and let $e$ denote the floor of $\frac{n}{2}$, which we write as $\lfloor\frac{n}{2}\rfloor$. Suppose $f\in Q$ is irreducible and the hypersurface $Q/(f)$ has an isolated singularity. Any nontrivial matrix factorization $(F^{0},F^{1},s^{0},s^{1})$ of $f$ satisfies $\operatorname{rank}(F^{0})\geq~{}2^{e}$.
We explain how Conjecture 1.1 is a weakening of [BGS87, Conjecture A] in Remark 2.2. The case of Conjecture 1.1 where $f$ is quadratic follows from Knörrer periodicity [Knö87]. Conjecture 1.1 was also recently proven to hold generically by Erman [Erm21]. Aside from these and several additional scattered cases, the conjecture remains open.
A graded version of Buchweitz-Greuel-Schreyer’s Conjecture implies a Horrocks-type splitting criterion for vector bundles of sufficiently small rank on projective hypersurfaces: see [BGS87, Conjecture B]. In this paper, we pose a quite different conjecture concerning the cohomology of sheaves on projective hypersurfaces, which is also motivated by [BGS87, Conjecture A]. Before stating it, we introduce some notation: given a proper scheme $Y$ over a field $k$ and a bounded complex $\mathcal{C}$ of coherent sheaves on $Y$, we set
$h^{j}(Y,\mathcal{C})\coloneqq\dim_{k}\mathbf{R}^{j}\Gamma(Y,\mathcal{C})\quad\text{and}\quad h(Y,\mathcal{C})\coloneqq\sum_{j\in\mathbb{Z}}h^{j}(Y,\mathcal{C}).$
Let $i\colon Y\hookrightarrow\mathbb{P}^{n}_{k}$ be a closed embedding. Given a bounded complex $\mathcal{C}$ of coherent sheaves on $Y$, we define
(1.2) $\rho(\mathcal{C})\coloneqq\sum_{r=0}^{n}h(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})\otimes\Omega_{\mathbb{P}^{n}_{k}}^{r}(r)).$
We caution the reader that $\rho(\mathcal{C})$ depends not just on the object $\mathcal{C}$ but also on the embedding $i\colon Y\hookrightarrow\mathbb{P}^{n}_{k}$.
Conjecture 1.3.
Let $k$ be an algebraically closed field and $i\colon X=V(f)\hookrightarrow\mathbb{P}^{n}_{k}$ a closed embedding of a smooth irreducible hypersurface of degree $d$ such that $a\coloneqq n+1-d\leq 0$ (i.e. such that $X$ is not Fano). For any nonzero object $\mathcal{C}\in\operatorname{D}^{\operatorname{b}}(X)$, we have $\rho(\mathcal{C})\geq 2^{e+1}$, where $e\coloneqq\lfloor\frac{n}{2}\rfloor$.
The lower bound in Conjecture 1.3 is sharp: see Example 4.7. Moreover, the statement in Conjecture 1.3 is false without the assumption that $a\leq 0$: see Remark 4.6. Conjecture 1.3 predicts that the total rank of the $E_{1}$ page of the Beilinson spectral sequence associated to $\mathcal{C}$ is at least $2^{e+1}$: see Remark 3.4 for details.
The connection between Conjectures 1.1 and 1.3 is illustrated by the following result:
Theorem 1.4.
Let $k$ be an algebraically closed field and $i\colon X=V(f)\hookrightarrow\mathbb{P}^{n}_{k}$ a closed embedding of a smooth irreducible hypersurface of degree $d$ such that $a\coloneqq n+1-d\leq 0$ (i.e. such that $X$ is not Fano).
  1. (1)
    Assume the graded version of Conjecture 1.1 holds for $f$. That is, assume all nontrivial graded matrix factorizations of $f$ (see Section 2.1 for the definition of a graded matrix factorization) satisfy the inequality in Conjecture 1.1. In this case, Conjecture 1.3 holds for $X$.
  2. (2)
    Assume $a=0$, i.e. $X$ is Calabi-Yau. If Conjecture 1.3 holds for $X$, then the graded version of Conjecture 1.1 holds for $f$.
See Theorem 3.3 for a more precise result concerning the relationship between ranks of matrix factorizations and sheaf cohomology, which we use to prove Theorem 1.4. Theorem 3.3 is an application of a result of Orlov relating graded singularity categories to derived categories of sheaves [Orl09, Theorem 2.5]. See also work of Pavlov [Pav20], which applies this same theorem of Orlov to give a cohomological formula for the Betti numbers of maximal Cohen-Macaulay modules over the coordinate ring of a Calabi-Yau variety in the case where the coordinate ring is Koszul.
Conventions
We index cohomologically. Given a complex $(C,d_{C})$ and $i\in\mathbb{Z}$, the shift $C[i]$ is given by $C[i]^{j}=C^{i+j}$ and $d_{C[i]}=(-1)^{i}d_{C}$. If $Q=\bigoplus_{i\in\mathbb{Z}}Q_{i}$ is a graded ring, and a $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ is a graded $Q$-module, then the $i^{\operatorname{th}}$ grading twist $M(i)$ of $M$ is given by $M(i)_{j}=M_{i+j}$. Given graded $Q$-modules $M$ and $N$, the set of degree 0 maps from $M$ to $N$ is denoted $\operatorname{Hom}_{R}(M,N)$, and we write the graded $R$-module given by the internal $\operatorname{Hom}$ from $M$ to $N$ as $\operatorname{\underline{Hom}}_{R}(M,N)$. The derived versions of these objects are denoted by $\mathbf{R}$$\operatorname{Hom}_{R}(M,N)$ and $\mathbf{R}$$\operatorname{\underline{Hom}}_{R}(M,N)$.
Setup 1.5.
We will use the following notation throughout the paper. Let $k$ be a field, $S\coloneqq k[x_{0},\dots,x_{n}]$ a standard graded polynomial ring, $f\in S$ a homogeneous and irreducible polynomial of degree $d$, $R\coloneqq S/(f)$, ${\mathfrak{m}}$ the homogeneous maximal ideal of $R$, $X$ the projective hypersurface $\operatorname{Proj}(R)$, and $i$ the closed embedding $X\hookrightarrow\mathbb{P}^{n}_{k}$. We assume $n\geq 1$ throughout (otherwise $X$ is empty). We write $a\coloneqq n+1-d$ and $e\coloneqq\lfloor\frac{n}{2}\rfloor$.
2. Background
Let $\operatorname{D}^{\operatorname{b}}_{\operatorname{gr}}(R)$ denote the bounded derived category of graded $S$-modules, $\operatorname{Perf}_{\operatorname{gr}}(R)\subseteq\operatorname{D}^{\operatorname{b}}_{\operatorname{gr}}(R)$ the subcategory of perfect complexes, and $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ the graded singularity category of $R$, i.e. the Verdier quotient $\operatorname{D}^{\operatorname{b}}_{\operatorname{gr}}(R)/\operatorname{Perf}_{\operatorname{gr}}(R)$. The graded singularity category can be interpreted in terms of graded matrix factorizations: let us recall some background on these objects.
2.1. Background on graded matrix factorizations
A graded matrix factorization of $f$ is a tuple $(F^{0},F^{1},s^{0},s^{1})$, where $F^{0}$, $F^{1}$ are finitely generated graded free $S$-modules, and
$s^{0}\colon F^{0}\to F^{1},\quad s^{1}\colon F^{1}(-d)\to F^{0}$
are homogeneous maps of degree 0 such that $s^{1}s^{0}=f\cdot\operatorname{id}_{F^{0}}$ and $s^{0}s^{1}=f\cdot\operatorname{id}_{F^{1}}$. We have $\operatorname{rank}(F^{0})=\operatorname{rank}(F^{1})$ [Eis80, Corollary 5.4]. A morphism of graded matrix factorizations $\alpha\colon(F^{0},F^{1},s^{0},s^{1})\to(G^{0},G^{1},t^{0},t^{1})$ is a pair of degree 0 maps $\alpha^{0}\colon F^{0}\to G^{0},\alpha^{1}\colon F^{1}\to G^{1}$ making the diagram
[…]
commute. A homotopy between such morphisms $\alpha,\beta$ is a pair of homogeneous maps $h^{0}\colon F^{0}\to G^{1}(-d)$, $h^{1}\colon F^{1}\to G^{0}$ satisfying the evident relations. The category $\operatorname{mf}_{\operatorname{gr}}(f)$ has objects given by graded matrix factorizations and morphisms as above, and the homotopy category $\operatorname{hmf}_{\operatorname{gr}}(f)$ of graded matrix factorizations has the same objects as $\operatorname{mf}_{\operatorname{gr}}(f)$ and morphisms given by modding out the maps in $\operatorname{mf}_{\operatorname{gr}}(f)$ by homotopy.
The category $\operatorname{hmf}_{\operatorname{gr}}(f)$ may be equipped canonically with the structure of a triangulated category; for instance, its shift functor sends $F=(F^{0},F^{1},s^{0},s^{1})$ to $F[1]\coloneqq(F^{1},F^{0}(d),-s^{1},-s^{0})$. Given $i\in\mathbb{Z}$, the $i^{\operatorname{th}}$ grading twist of $F=(F^{0},F^{1},s^{0},s^{1})$ is $F(i)=(F^{0}(i),F^{1}(i),s^{0},s^{1})$. Observe that $F[2]=F(d).$
A finitely generated $R$-module $M$ is called maximal Cohen-Macaulay (MCM) if $0pt(M)=\dim(M)$. Let $\underline{\operatorname{MCM}}_{\operatorname{gr}}(R)$ denote the stable category of graded MCM $R$-modules, i.e. the category with objects given by graded MCM $R$-modules and morphisms given by $R$-linear maps modulo those that factor through a projective module. The category $\underline{\operatorname{MCM}}_{\operatorname{gr}}(R)$ may be equipped with a triangulated structure such that the shift functor sends an MCM module $M$ to its first syzygy. It follows from (a graded version of) a theorem of Eisenbud [Eis80, Theorem 6.3] that there is an equivalence $\operatorname{hmf}_{\operatorname{gr}}(f)\xrightarrow{\simeq}\underline{\operatorname{MCM}}_{\operatorname{gr}}(R)$ of triangulated categories given by $(F^{0},F^{1},s^{0},s^{1})\mapsto\operatorname{coker}(s^{0})$. Moreover, it follows from (a graded version of) a theorem of Buchweitz [Buc21, Theorem 4.4.1(2)] that there is an equivalence $\underline{\operatorname{MCM}}_{\operatorname{gr}}(R)\xrightarrow{\simeq}\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ of triangulated categories that sends a module to itself, considered as a complex concentrated in cohomological degree 0. Composing these two functors yields an equivalence of triangulated categories
(2.1) $\operatorname{coker}\colon\operatorname{hmf}_{\operatorname{gr}}(f)\xrightarrow{\simeq}\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R).$
Notice that $\operatorname{coker}(F(i))=\operatorname{coker}(F)(i)$ for all $F\in\operatorname{hmf}_{\operatorname{gr}}(f)$ and $i\in\mathbb{Z}$.
Remark 2.2.
Let us explain the relationship between Buchweitz-Greuel-Schreyer’s Conjecture [BGS87, Conjecture A] and Conjecture 1.1. The former states that, in the setting of Conjecture 1.1, if $M$ is a nonfree MCM $Q/(f)$-module, then $\operatorname{rank}(M)\geq 2^{e-1}$. [^1] Buchweitz-Greuel-Schreyer assume in [BGS87, Conjecture A] that $M$ has no free summands, but there is no loss in assuming $M$ is simply nonfree. Also, while [BGS87, Conjecture A] does not contain an isolated singularity assumption, Buchweitz-Greuel-Schreyer remark directly below [BGS87, Conjecture A] that the conjecture reduces immediately to the isolated singularity case. Given a nontrivial matrix factorization $(F^{0},F^{1},s^{0},s^{1})$ of $f$, the $Q/(f)$-modules $\operatorname{coker}(s^{0})$ and $\operatorname{coker}(s^{1})$ are both nonfree and maximal Cohen-Macaulay; since $\operatorname{rank}(F^{0})=\operatorname{rank}(\operatorname{coker}(s^{0}))+\operatorname{rank}(\operatorname{coker}(s^{1}))$, [BGS87, Conjecture A] implies Conjecture 1.1.
2.2. Orlov’s Theorem
The following is (a special case of) a celebrated theorem of Orlov:
Theorem 2.3 ([Orl09] Theorem 2.5).
There are functors
$\Phi_{i}\colon\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)\to\operatorname{D}^{\operatorname{b}}(X)\quad\text{and}\quad\Psi_{i}\colon\operatorname{D}^{\operatorname{b}}(X)\to\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$
for all $i\in\mathbb{Z}$ satisfying the following:
  1. (1)
    If $a>0$ (i.e. if $X$ is Fano), then $\Phi_{i}$ is fully faithful, and there is a semiorthogonal decomposition $\operatorname{D}^{\operatorname{b}}(X)=\langle\mathcal{O}(-i-a+1),\dots,\mathcal{O}(-i-1),\mathcal{O}(-i),\Phi_{i}(\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R))\rangle.$
  2. (2)
    If $a<0$ (i.e. if $X$ is general type), then $\Psi_{i}$ is fully faithful, and there is a semiorthogonal decomposition $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)=\langle k(-i),k(-i-1)\dots,k(-i+a+1),\Psi_{i}(\operatorname{D}^{\operatorname{b}}(X))\rangle.$
  3. (3)
    If $a=0$ (i.e. if $X$ is Calabi-Yau), then $\Phi_{i}$ and $\Psi_{i}$ are inverse equivalences.
Let us recall the formulas for the functors $\Phi_{i}$ and $\Psi_{i}$:
Definition of the functors $\Phi_{i}$
We follow Burke-Stevenson’s exposition in [BS15] of the proof of Orlov’s Theorem. The functor $\Phi_{i}$ factors as a composition $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)\xrightarrow{\varphi_{i}}\operatorname{D}^{\operatorname{b}}_{\operatorname{gr}}(R)\to\operatorname{D}^{\operatorname{b}}(X),$ where the second map is given by sheafification. Before defining $\varphi_{i}$, we fix some notation and recall some background. Given a complex $G$ of graded free $R$-modules, we let $G_{\prec i}$ be the subcomplex of $G$ given by free summands of the form $R(i)$ with $i>0$, and we define $G_{\succeq i}\coloneqq G/G_{\prec i}$. We recall that any complex $C$ of finitely generated graded $R$-modules such that $H^{j}(C)=0$ for $j\gg 0$ admits a minimal free resolution, i.e. a complex $(F,d_{F})$ of graded free $R$-modules such that $F^{j}=0$ for $j\gg 0$, $d_{F}(F)\subseteq{\mathfrak{m}}F$, and there is a quasi-isomorphism $F\xrightarrow{\simeq}C$. Such a resolution exists and is unique by [Rob98, Proposition 4.4.1].
Suppose $M$ is a bounded complex of finitely generated graded $R$-modules, i.e. an object in $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$: the object $\varphi_{i}(M)$ is defined as follows. Let $F$ be the minimal $R$-free resolution of $M$, and let $F^{\prime}$ be the minimal $R$-free resolution of the complex $\operatorname{\underline{Hom}}_{R}(F_{\succeq i},R)$. Observe that $\operatorname{\underline{Hom}}_{R}(F_{\succeq i},R)$ is a complex of graded free $R$-modules, and it has bounded cohomology since $R$ is Gorenstein; also, $\operatorname{\underline{Hom}}_{R}(F_{\succeq i},R)$ is typically not bounded above, and so it usually does not coincide with $F^{\prime}$. Finally, we set $\varphi_{i}(M)\coloneqq\operatorname{\underline{Hom}}_{R}(F^{\prime},R)_{\prec i}$.
Example 2.4.
Suppose $M$ is a bounded complex of finitely generated graded free $R$-modules; we observe that $\varphi_{i}(M)=0$ for all $i$. Indeed, we have $F=M$, and $F^{\prime}=\operatorname{\underline{Hom}}_{R}(M_{\succeq i},R)$; thus, $\varphi_{i}(M)=(M_{\succeq i})_{\prec i}=0$.
Example 2.5.
[…]
Definition of the functors $\Psi_{i}$
The functor $\Psi_{i}$ factors as the composition $\operatorname{D}^{\operatorname{b}}(X)\xrightarrow{\psi_{i}}\operatorname{D}^{\operatorname{b}}_{\operatorname{gr}}(R)\to\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R),$ where $\psi_{i}(\mathcal{F})=\bigoplus_{j\geq i-a}\mathbf{R}\Gamma(X,\mathcal{F}(j))$, and the second map is the canonical one [Orl09, Remark 2.6].
Proposition 2.6.
In the context of Theorem 2.3, the functor $\Phi_{i}$ is the left adjoint of the functor $\Psi_{i+a}$.
Proof.
Given $i\in\mathbb{Z}$, let $\operatorname{D}^{\operatorname{b}}(R)_{\geq i}$ denote the subcategory of $\operatorname{D}^{\operatorname{b}}(R)$ given by objects whose cohomology is concentrated in internal degrees at least $i$. It follows from the proof of [BS15, Lemma 5.4] that $\varphi_{i}$ takes values in $\operatorname{D}^{\operatorname{b}}(R)_{\geq i}$. The functor $\Phi_{i}$ therefore factors as the composition $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)\xrightarrow{\varphi_{i}}\operatorname{D}^{\operatorname{b}}(R)_{\geq i}\xrightarrow{\pi_{i}}\operatorname{D}^{\operatorname{b}}(X),$ where $\pi_{i}$ is given by sheafification. On the other hand, the functor $\Psi_{i+a}$ may be expressed as the composition $\operatorname{D}^{\operatorname{b}}(X)\xrightarrow{\psi_{i+a}}\operatorname{D}^{\operatorname{b}}(R)_{\geq i}\xrightarrow{\pi^{\prime}_{i}}\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R),$ where $\pi^{\prime}_{i}$ is the canonical surjection. The map $\varphi_{i}$ is the left adjoint of $\pi^{\prime}_{i}$ [BS15, Proposition 5.8], and $\pi_{i}$ is the left adjoint of $\psi_{i+a}$; the result follows. ∎
2.3. Graded Auslander duality
Let $A=\bigoplus_{i\geq 0}A_{i}$ be a commutative, graded Gorenstein ring of dimension $t$ such that $A_{0}=k$, and assume $A$ has an isolated singularity. There is an isomorphism $\mathbf{R}{\operatorname{\underline{Hom}}}_{A}(k,A)\cong k(b)[-t]$ in $\operatorname{D}^{\operatorname{b}}_{\operatorname{gr}}(A)$ for some $b,t\in\mathbb{Z}$. For example, our graded hypersurface ring $R$ is Gorenstein, and in this case $b$ is equal to $a=n+1-d$, and $t$ is equal to $n$. The following result is a version of Auslander duality [Aus78] for the graded singularity category $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(A)$ (see also [Mur13, (3.7)]). Given a graded $k$-vector space $V$, let $V^{*}\coloneqq\operatorname{\underline{Hom}}_{k}(V,k)$ denote its $k$-dual.
Proposition 2.7 ([KMVdB11] Section 4.3).
For any $M,N\in\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(A)$, there is a natural isomorphism
$\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(A)}(M,N)^{*}\cong\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(A)}(N,M(-b)[t-1]).$
That is, the automorphism $M\mapsto M(-b)[t-1]$ of $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(A)$ is a Serre functor.
3. Proof of Theorem 1.4
We will need the following technical result.
Proposition 3.1.
Assume $a\leq 0$. Let $\ell\in\mathbb{Z}$, and consider $k(\ell)$ as a complex of graded $R$-modules concentrated in cohomological degree 0. Write $\ell=qd-r$, where $0\leq r<d$. Let $\mathcal{T}_{\mathbb{P}^{n}_{k}}$ denote the tangent bundle on $\mathbb{P}^{n}_{k}$. There is an isomorphism $\Phi_{0}(k(\ell))\cong i^{*}(\bigwedge^{r+a}\mathcal{T}_{\mathbb{P}^{n}_{k}})(-r-a)[2q+n-r-a-1]$ in $\operatorname{D}^{\operatorname{b}}(X)$.
Proof.
We have $F[2]=F(d)$ for all $F\in\operatorname{hmf}_{\operatorname{gr}}(f)$, and the equivalence in (2.1) commutes with grading twists and cohomological shifts; it follows that $k(\ell)[2j]\cong k(\ell+jd)$ in $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ for all $j\in\mathbb{Z}$. Since the functor $\Phi_{0}$ commutes with cohomological shifts, we may assume $q=0$, i.e. $-d<\ell\leq 0$. We therefore must show $\Phi_{0}(k(\ell))\cong i^{*}(\bigwedge^{a-\ell}\mathcal{T}_{\mathbb{P}^{n}_{k}}(\ell-a))[n+\ell-a-1]$.
Let $F$ be the minimal $R$-free resolution of $k$. The resolution $F$ is given by applying the Shamash construction [EP16, Construction 4.1.3] to the Koszul complex
$K\coloneqq\left[S(-n-1)\to S(-n)^{n+1}\to\cdots\to S(-1)^{n+1}\to S\right]$
on the variables $x_{0},\dots,x_{n}\in S$. The underlying module of $F$ is thus $K\otimes_{S}D\otimes_{S}R$, where $D$ is a divided power algebra on a single variable $t$ of cohomological degree $-2$ and internal degree $d$. We therefore have:
$F^{m}=\bigoplus_{-s-2j=m,j\geq 0}S(-s)^{\binom{n+1}{s}}\otimes_{S}S(-jd)\cdot t^{(j)}\otimes_{S}R=\bigoplus_{-s-2j=m,j\geq 0}R(-s-jd)^{\binom{n+1}{s}}.$
The first few terms of $F$ are
$\left[\cdots\to R(-3)^{\binom{n+1}{3}}\oplus R(-d-1)^{n+1}\to R(-2)^{\binom{n+1}{2}}\oplus R(-d)\to R(-1)^{n+1}\to R\right].$
We have $F(\ell)_{\succeq 0}=F(\ell)$ (using the truncation notation from the description of the functor $\Phi_{0}$ in Section 2.1). Since $\mathbf{R}$$\operatorname{\underline{Hom}}_{R}(k,R)\cong k(a)[-n]$, we have an isomorphism
$\operatorname{\underline{Hom}}_{R}(F(\ell)_{\succeq 0},R)=\operatorname{\underline{Hom}}_{R}(F(\ell),R)\cong k(a-\ell)[-n]$
in $\operatorname{D}^{\operatorname{b}}(R)$, and so the minimal free resolution of $\operatorname{\underline{Hom}}_{R}(F(\ell)_{\succeq 0},R)$ is $F^{\prime}\coloneqq F(a-\ell)[-n]$. We now compute:
$\varphi_{0}(k(\ell))$ $=\operatorname{\underline{Hom}}_{R}(F^{\prime},R)_{\prec 0}$
$=\operatorname{\underline{Hom}}_{R}(F,R)(\ell-a)[n]_{\prec 0}$
$=\left[R(\ell-a)\to R(\ell-a+1)^{n+1}\to R(\ell-a+2)^{\binom{n+1}{2}}\oplus R(\ell-a+d)\to\cdots\right]_{\prec 0},$
where $R(\ell-a)$ is in cohomological degree $-n$. Observe that
$\operatorname{\underline{Hom}}_{R}(F,R)(\ell-a)[n]\cong\text{$\mathbf{R}$$\operatorname{\underline{Hom}}_{R}(k,R)(\ell-a)[n]$}\cong k(\ell);$
in particular, the sheafification of $\operatorname{\underline{Hom}}_{R}(F,R)(\ell-a)[n]$ is the zero object in $\operatorname{D}^{\operatorname{b}}(X)$. It follows that $\Phi_{0}(k(\ell))$ is isomorphic to the sheafification of the complex
$\left[R(\ell-a)\to R(\ell-a+1)^{n+1}\to R(\ell-a+2)^{\binom{n+1}{2}}\oplus R(\ell-a+d)\to\cdots\right]_{\succeq 0},\$
where $R(\ell-a)$ is in cohomological degree $-n+1$. Notice that $\ell-a+d>0$ (this is where we use the assumption $a\leq 0$). Thus, if $\ell>a$, then $\Phi_{0}(k(\ell))=0$; otherwise, $\Phi_{0}(k(\ell))$ is isomorphic to
$\left[\mathcal{O}_{X}(\ell-a)\to\mathcal{O}_{X}(\ell-a+1)^{n+1}\to\mathcal{O}_{X}(\ell-a+2)^{\binom{n+1}{2}}\to\cdots\to\mathcal{O}_{X}^{\binom{n+1}{a-\ell}}\right],$
a brutally truncated Koszul complex on $x_{0},\dots,x_{n}$, where $\mathcal{O}_{X}(\ell-a)$ is in cohomological degree $-n+1$. This complex is a locally free resolution of the sheaf $i^{*}\operatorname{ker}\left(\mathcal{O}_{\mathbb{P}^{n}_{k}}(1)^{\binom{n+1}{a-\ell+1}}\to\mathcal{O}_{\mathbb{P}^{n}_{k}}(2)^{\binom{n+1}{a-\ell+2}}\right)\cong i^{*}(\bigwedge^{a-\ell}\mathcal{T}_{\mathbb{P}^{n}_{k}})(\ell-a)$ in cohomological degree $a-\ell-n+1$, where the isomorphism follows from [Eis95, Theorem 17.16]. ∎
Remark 3.2.
In the setting of Proposition 3.1, assume $X$ is smooth. Combining a result of Tu [Tu14, Theorem 1.4] and the equivalence (2.1), one sees that the category $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ is generated by the objects $k,\dots,k(d-1)$. Proposition 3.1 therefore implies that the objects $i^{*}(\bigwedge^{r}\mathcal{T}_{\mathbb{P}^{n}_{k}})(-r)$ with $0\leq r<d$ generate $\operatorname{D}^{\operatorname{b}}(X)$.
A graded matrix factorization $(F^{0},F^{1},s^{0},s^{1})$ is called reduced if the maps $s^{0}$ and $s^{1}$ are minimal, i.e. $s^{0}\otimes k=s^{1}\otimes k=0$.
Theorem 3.3.
Assume $X$ is smooth and $a\leq 0$, and let $\mathcal{C}\in\operatorname{D}^{\operatorname{b}}(X)$. Let $F=(F^{0},F^{1},s^{0},s^{1})$ be a reduced graded matrix factorization of $f$ such that $\Psi_{a}(\mathcal{C})\cong\operatorname{coker}(F)$, where $\Psi_{a}$ is as in Theorem 2.3. For $i\in\{0,1\}$ and $j\in\mathbb{Z}$, write $b^{i}_{j}\coloneqq\dim_{k}\operatorname{Hom}_{S}(F^{i},k(-j))$, the number of degree $j$ generators of $F^{i}$.
  1. (1)
    Fix $i\in\{0,1\}$ and $j\in\mathbb{Z}$, and write $a-j=qd-r$ for $0\leq r<d$. We have:
    $b^{i}_{j}=h^{r+a-2q-i+1}\left(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})\otimes\Omega^{r+a}_{\mathbb{P}^{n}_{k}}(r+a)\right).$
  2. (2)
    Additionally, $\operatorname{rank}(F)=\rho(\mathcal{C})$, where $\rho$ is as defined in (1.2).
Proof.
Write $M\coloneqq\operatorname{coker}(F)$. Since $F$ is reduced, the minimal free resolution of $M$ has the form:
$\cdots\xrightarrow{s^{1}}F^{0}(-d)\xrightarrow{s^{0}}F^{1}(-d)\xrightarrow{s^{1}}F^{0}\xrightarrow{s^{0}}F^{1}\to M.$
Thus, the graded Betti number of $M$ in cohomological degree 0 (resp. $-1$) and internal degree $j$ is $b^{1}_{j}$ (resp. $b^{0}_{j})$. We now compute:
$b^{i}_{j}$ $=\dim_{k}\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)}\left(\Psi_{a}(\mathcal{C}),k(-j)[i-1]\right)$
$=\dim_{k}\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)}\left(k(a-j),\Psi_{a}(\mathcal{C})[n-i]\right)$
$=\dim_{k}\operatorname{Ext}^{n-i}_{X}\left(\Phi_{0}(k(a-j)),\mathcal{C}\right)$
$=\dim_{k}\operatorname{Ext}_{X}^{r+a+1-i-2q}\left(i^{*}(\bigwedge^{r+a}\mathcal{T}_{\mathbb{P}^{n}_{k}})(-r-a),\mathcal{C}\right)$
$=h^{r+a+1-i-2q}\left(X,\mathcal{C}\otimes i^{*}(\Omega^{r+a}_{\mathbb{P}^{n}_{k}})(r+a)\right)$
$=h^{r+a+1-i-2q}\left(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})\otimes\Omega^{r+a}_{\mathbb{P}^{n}_{k}}(r+a)\right),$
where the first equality follows from [Pav21, Lemma 3.1] and the identification of the shift $M[1]$ in $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ with the first syzygy of $M$, the second from Proposition 2.7, the third from Proposition 2.6, the fourth from Theorem 3.1, the fifth from a straightforward calculation, and the sixth from the projection formula. This proves (1). For (2), we have $\operatorname{rank}(F)=\operatorname{rank}(F^{0})+\operatorname{rank}(F^{1})=\sum_{j\in\mathbb{Z}}(b^{0}_{j}+b^{1}_{j})~{}=\rho(\mathcal{C}).$
Proof of Theorem 1.4.
Assume the graded version of Conjecture 1.1 holds for $f$. Let $\mathcal{C}\in\operatorname{D}^{\operatorname{b}}(X)$. Choose $F=(F^{0},F^{1},s^{0},s^{1})\in\operatorname{mf}_{\operatorname{gr}}(f)$ such that $\operatorname{coker}(F)\cong~{}\Psi_{a}(\mathcal{C})$. We have $2^{e+1}\leq 2\cdot\operatorname{rank}(F^{0})=\rho(\mathcal{C}),$ where the equality follows from Theorem 3.3(2). This proves Part (1). Assume Conjecture 1.3 holds, let $F$ be a nontrivial graded matrix factorization of $f$, and let $\mathcal{C}\coloneqq(\Phi_{0}\circ\operatorname{coker})(F)$. By Theorem 2.3(3) and Theorem 3.3(2), we have $2\cdot\operatorname{rank}(F^{0})=\rho(\mathcal{C})\geq 2^{e+1}.$
Remark 3.4.
Given any object $\mathcal{C}\in\operatorname{D}^{\operatorname{b}}(X)$, there is a second-quadrant spectral sequence
$E_{1}^{p,q}=H^{q}\left(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})\otimes\Omega_{\mathbb{P}^{n}_{k}}^{-p}(-p)\right)\otimes_{k}\mathcal{O}(p)\Rightarrow\mathcal{H}^{p+q}(i_{*}(\mathcal{C}))$
called the Beilinson spectral sequence [Beĭ78]. Notice that the total rank of the $E_{1}$ page of this spectral sequence is exactly $\rho(\mathcal{C})$.
4. Examples
We first note that the invariant $\rho$ enjoys a duality property:
Proposition 4.1.
For any bounded complex $\mathcal{C}$ of locally free $\mathcal{O}_{X}$-modules, we have $\rho(\mathcal{C})=\rho(\mathcal{C}^{\vee}(1-a))$, where $\mathcal{C}^{\vee}\coloneqq\mathcal{H}om_{X}(\mathcal{C},\mathcal{O}_{X})$.
Proof.
Noting that the canonical bundle of $X$ is $\mathcal{O}_{X}(-a)$, we have:
$\rho(\mathcal{C})$ $\coloneqq\sum_{r=0}^{n}h\left(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})\otimes\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\right)$
$=\sum_{r=0}^{n}h\left(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})^{\vee}\otimes\bigwedge^{r}\mathcal{T}_{\mathbb{P}_{k}^{n}}(-r-a)\right)$
$=\sum_{r=0}^{n}h\left(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})^{\vee}\otimes\Omega_{\mathbb{P}^{n}_{k}}^{n-r}(n+1-r-a)\right)$
$=\sum_{r=0}^{n}h\left(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})^{\vee}(1-a)\otimes\Omega_{\mathbb{P}^{n}_{k}}^{r}(r)\right),$
where the second equality follows from Serre duality, the third from the isomorphism $\bigwedge^{r}\mathcal{T}_{\mathbb{P}_{k}^{n}}\cong\Omega_{\mathbb{P}^{n}_{k}}^{n-r}(n+1)$ for all $0\leq r\leq n$ (see, for instance, the proof of [Eis95, Theorem 17.16]), and the last from reindexing. ∎
We next observe that Conjecture 1.3 holds for points and for the structure sheaf $\mathcal{O}_{X}$.
Example 4.2.
Let $\mathcal{F}$ be the structure sheaf of a point on $X$. We evidently have
$\rho(\mathcal{F})=\sum_{r=0}^{n}\operatorname{rank}(\Omega^{r}_{\mathbb{P}^{n}_{k}})=2^{n+1}-1>2^{e+1}.$
Example 4.3.
By [OSS11] Chapter 1, Section 1.1], for all $0\leq p,q\leq n$ and $\ell\in\mathbb{Z}$, we have:
(4.4) $h^{q}(\mathbb{P}^{n}_{k},\Omega^{p}_{\mathbb{P}^{n}_{k}}(\ell))=\begin{cases}\binom{\ell+n-p}{\ell}\binom{\ell-1}{p},&q=0,\text{ }\ell>p;\\ 1,&\ell=0,\text{ }q=p;\\ \binom{p-\ell}{-\ell}\binom{-\ell-1}{n-p},&q=n,\text{ }\ell<p-n;\\ 0,&\operatorname{otherwise.}\end{cases}$
For each $0\leq r\leq n$, there is a short exact sequence
(4.5) $0\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r-d)\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to\mathcal{O}_{X}\otimes\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to 0$
Assume now that $a\leq 0$. By (4.4), we have:
  1. (1)
    $h^{i}\left(\mathbb{P}_{k}^{n},\Omega_{\mathbb{P}_{k}^{n}}^{r}(r-d)\right)=\begin{cases}\binom{d}{d-r}\binom{d-r-1}{n-r},&i=n;\\ 0,&i\neq n.\end{cases}$
  2. (2)
    $h^{i}\left(\mathbb{P}_{k}^{n},\Omega_{\mathbb{P}_{k}^{n}}^{r}(r)\right)=\begin{cases}1,&i=r=0;\\ 0,&\text{else.}\end{cases}$
The short exact sequence (4.5) therefore implies:
$\rho(\mathcal{O}_{X})=1+\sum_{r=0}^{n}\binom{d}{d-r}\binom{d-r-1}{n-r}\geq 1+\sum_{r=0}^{n}\binom{n+1}{n+1-r}=2^{n+1}>2^{e+1}.$
Remark 4.6.
The assumption in Conjecture 1.3 that $a\leq 0$ is necessary. For instance, suppose $n\geq 2$ and that $f$ is a linear form, so that $a=n>0$. For all $1-n\leq j\leq 0$, the calculation (4.4) and the short exact sequence (4.5) imply that
$\rho(\mathcal{O}_{X}(j))=h^{-j}(\mathbb{P}_{k}^{n},\mathcal{O}_{X}\otimes\Omega^{-j}_{\mathbb{P}^{n}_{k}})+h^{-j}(\mathbb{P}_{k}^{n},\mathcal{O}_{X}\otimes\Omega^{1-j}_{\mathbb{P}^{n}_{k}}(1))=h^{-j}(\mathbb{P}_{k}^{n},\Omega^{-j}_{\mathbb{P}^{n}_{k}})+h^{1-j}(\mathbb{P}_{k}^{n},\Omega^{1-j}_{\mathbb{P}^{n}_{k}})=2<2^{e+1}.$
Thus, each of the line bundles $\mathcal{O}_{X}(1-n),\dots,\mathcal{O}_{X}$ violates the inequality in Conjecture 1.3. Additionally, if $n=d=2$, then $\rho(\mathcal{O}_{X})=2<2^{e+1}=4$.
Finally, we consider a family of examples where the bound in Conjecture 1.3 is achieved.
Example 4.7.
Assume $k$ contains a square root $i$ of $-1$, and suppose $f=x_{0}^{n}+\cdots+x_{n}^{n}$, where $n$ is even. In this case, $e=\frac{n}{2}$. We have graded matrix factorizations
$F=\left(S(-e),S,x_{0}^{e},x_{0}^{e}\right)\in\operatorname{mf}_{\operatorname{gr}}(x_{0}^{n}),\quad F^{\prime}=\left(S(-e),S,x_{0}^{e}+ix_{1}^{e},x_{0}^{e}-ix_{1}^{e}\right)\in\operatorname{mf}_{\operatorname{gr}}(x_{0}^{n}+x_{1}^{n}).$
We may construct a graded matrix factorization of $f$ whose rank meets the bound in Conjecture 1.1 by taking tensor products of $F$ and $F^{\prime}$; we refer the reader to [BD20, Section 4] for the definition of the tensor product of graded matrix factorizations. In more detail: the tensor product of one copy of $F$ and $e$ copies of $F^{\prime}$ gives a reduced graded matrix factorization $G=(G^{0},G^{1},s^{0},s^{1})$ of $f$, where $G^{0}=S(-e)^{2^{e}}$, and $G^{1}=S^{2^{e}}$. Let $M\coloneqq\operatorname{coker}(s^{0})$; since $M$ is MCM and generated in degree 0, $\Phi_{0}(M)$ is the vector bundle $\widetilde{M}$ (Example 2.5). Since $a=0$, we conclude from Theorem 2.3(3) that $\Psi_{0}(\widetilde{M})=M$. By Theorem 3.3(2), we therefore have $\rho(\widetilde{M})=2^{e+1}$, i.e. $\widetilde{M}$ achieves the bound in Conjecture 1.3.

### v2

@@
 Let $k$ be a field, $S\coloneqq k[x_{0},\dots,x_{n}]$ a standard graded polynomial ring, $f\in S$ a homogeneous and irreducible polynomial of degree $d$, $R\coloneqq S/(f)$, $\mathfrak{m}$} the homogeneous maximal ideal of $R$, $X$ the projective hypersurface $\operatorname{Proj}(R)$, and $i$ the closed embedding $X\hookrightarrow\mathbb{P}^{n}_{k}$.
-We assume $n\geq 1$ throughout (otherwise $X$ is empty).
+We assume $n\geq 1$ throughout.
 We write $a\coloneqq n+1-d$ and $e\coloneqq\lfloor\frac{n}{2}\rfloor$.
@@
 A morphism of graded matrix factorizations $\alpha\colon(F^{0},F^{1},s^{0},s^{1})\to(G^{0},G^{1},t^{0},t^{1})$ is a pair of degree 0 maps $\alpha^{0}\colon F^{0}\to G^{0},\alpha^{1}\colon F^{1}\to G^{1}$ making the diagram
-[…]
+${F^{1}(-d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s^{1}}$$\scriptstyle{\alpha^{1}}$${F^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s^{0}}$$\scriptstyle{\alpha^{0}}$${F^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha^{1}}$${G^{1}(-d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{t^{1}}$${G^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{t^{0}}$${G^{1}}$
 commute.
@@
 Example 2.5.
-[…]
+Suppose $M$ is an MCM module generated in degree 0; we will discuss explicit examples of such modules in Example 4.7 below.
+We now show $\Phi_{0}(M)=\widetilde{M}$.
+Let $F$ be the minimal free resolution of $M$, which has the form $F=\left[P\xleftarrow{\alpha}Q\xleftarrow{\beta}P(-d)\xleftarrow{\alpha}Q(-d)\xleftarrow{\beta}\cdots\right]$ for some graded free $R$-modules $P$ and $Q$ and maps $\alpha$ and $\beta$, where $P$ is a direct sum of copies of $R$.
+Given a graded $R$-module NN, write $N^{\vee}\coloneqq\operatorname{\underline{Hom}}_{R}(N,R)$.
+We have $F_{\succeq 0}=F$, and $\operatorname{\underline{Hom}}_{R}(F,R)=\left[P^{\vee}\xrightarrow{\alpha^{\vee}}Q^{\vee}\xrightarrow{\beta^{\vee}}P^{\vee}(d)\xrightarrow{\alpha^{\vee}}Q^{\vee}(d)\xrightarrow{\beta^{\vee}}\cdots\right]$.
+The minimal free resolution of $\operatorname{\underline{Hom}}_{R}(F,R)$ is the complex $F^{\prime}=\left[Q^{\vee}(-d)\xleftarrow{\alpha^{\vee}}P^{\vee}(-d)\xleftarrow{\beta^{\vee}}Q^{\vee}(-2d)\xleftarrow{\alpha^{\vee}}P^{\vee}(-2d)\xleftarrow{\beta^{\vee}}\cdots\right]$.
+Dualizing $F^{\prime}$ gives the complex $\left[Q(d)\xrightarrow{\alpha}P(d)\xrightarrow{\beta}Q(2d)\xrightarrow{\alpha}P(2d)\xrightarrow{\beta}\cdots\right]$, whose terms are direct sums of copies of $R(j)$ for various $j>0$.
+We conclude that $\varphi_{0}(M)=\operatorname{\underline{Hom}}_{R}(F^{\prime},R)$, which is quasi-isomorphic to $M$.
+Thus, $\Phi_{0}(M)=\widetilde{M}$.
 Definition of the functors $\Psi_{i}$
@@
 There is an isomorphism $\mathbf{R}{\operatorname{\underline{Hom}}}_{A}(k,A)\cong k(b)[-t]$ in $\operatorname{D}^{\operatorname{b}}_{\operatorname{gr}}(A)$ for some $b,t\in\mathbb{Z}$.
-For example, our graded hypersurface ring $R$ is Gorenstein, and in this case $b$ is equal to $a=n+1-d$, and $t$ is equal to $n$.
+For example, our graded hypersurface ring $R$ is Gorenstein, and in this case $b$ is equal to a=n+1da=n+1-d, and $t$ is equal to $n$.
 The following result is a version of Auslander duality [Aus78] for the graded singularity category $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(A)$ (see also [Mur13, (3.7)]).
@@
 Since the functor $\Phi_{0}$ commutes with cohomological shifts, we may assume $q=0$, i.e.
-$-d<\ell\leq 0$.
+$d<\ell\leq 0$.
 We therefore must show $\Phi_{0}(k(\ell))\cong i^{*}(\bigwedge^{a-\ell}\mathcal{T}_{\mathbb{P}^{n}_{k}}(\ell-a))[n+\ell-a-1]$.
@@
 on the variables $x_{0},\dots,x_{n}\in S$.
-The underlying module of $F$ is thus $K\otimes_{S}D\otimes_{S}R$, where $D$ is a divided power algebra on a single variable $t$ of cohomological degree $-2$ and internal degree $d$.
+The underlying module of $F$ is thus $K\otimes_{S}D\otimes_{S}R$, where $D$ is a divided power algebra on a single variable $t$ of cohomological degree 2-2 and internal degree $d$.
 We therefore have:
@@
 $=\left[R(\ell-a)\to R(\ell-a+1)^{n+1}\to R(\ell-a+2)^{\binom{n+1}{2}}\oplus R(\ell-a+d)\to\cdots\right]_{\prec 0},$
-where $R(\ell-a)$ is in cohomological degree $-n$. Observe that
+where $R(\ell-a)$ is in cohomological degree n-n. Observe that
 $\operatorname{\underline{Hom}}_{R}(F,R)(\ell-a)[n]\cong\text{$\mathbf{R}$$\operatorname{\underline{Hom}}_{R}(k,R)(\ell-a)[n]$}\cong k(\ell);$
@@
 It follows that $\Phi_{0}(k(\ell))$ is isomorphic to the sheafification of the complex
-$\left[R(\ell-a)\to R(\ell-a+1)^{n+1}\to R(\ell-a+2)^{\binom{n+1}{2}}\oplus R(\ell-a+d)\to\cdots\right]_{\succeq 0},\$
-where $R(\ell-a)$ is in cohomological degree $-n+1$.
+$\left[R(\ell-a)\to R(\ell-a+1)^{n+1}\to R(\ell-a+2)^{\binom{n+1}{2}}\oplus R(\ell-a+d)\to\cdots\right]_{\succeq 0},\\$
+where $R(\ell-a)$ is in cohomological degree n+1-n+1.
 Notice that $\ell-a+d>0$ (this is where we use the assumption $a\leq 0$).
@@
 $\left[\mathcal{O}_{X}(\ell-a)\to\mathcal{O}_{X}(\ell-a+1)^{n+1}\to\mathcal{O}_{X}(\ell-a+2)^{\binom{n+1}{2}}\to\cdots\to\mathcal{O}_{X}^{\binom{n+1}{a-\ell}}\right$],
-a brutally truncated Koszul complex on $x_{0},\dots,x_{n}$, where $\mathcal{O}_{X}(\ell-a)$ is in cohomological degree $-n+1$.
+a brutally truncated Koszul complex on $x_{0},\dots,x_{n}$, where $\mathcal{O}_{X}(\ell-a)$ is in cohomological degree n+1-n+1.
 This complex is a locally free resolution of the sheaf $i^{*}\operatorname{ker}\left(\mathcal{O}_{\mathbb{P}^{n}_{k}}(1)^{\binom{n+1}{a-\ell+1}}\to\mathcal{O}_{\mathbb{P}^{n}_{k}}(2)^{\binom{n+1}{a-\ell+2}}\right)\cong i^{*}(\bigwedge^{a-\ell}\mathcal{T}_{\mathbb{P}^{n}_{k}})(\ell-a)$ in cohomological degree $a-\ell-n+1$, where the isomorphism follows from [Eis95, Theorem 17.16].
@@
   1. (1)
-    Fix $i\in\{0,1\}$ and $j\in\mathbb{Z}$, and write $a-j=qd-r$ for $0\leq r<d$. We have:
+    Fix $i\in\{0,1\}$ and $j\in\mathbb{Z}$, and write aj=qdra-j=qd-r for $0\leq r<d$. We have:
     $b^{i}_{j}=h^{r+a-2q-i+1}\left(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})\otimes\Omega^{r+a}_{\mathbb{P}^{n}_{k}}(r+a)\right).$
@@
 $\cdots\xrightarrow{s^{1}}F^{0}(-d)\xrightarrow{s^{0}}F^{1}(-d)\xrightarrow{s^{1}}F^{0}\xrightarrow{s^{0}}F^{1}\to M.$
-Thus, the graded Betti number of $M$ in cohomological degree 0 (resp. $-1$) and internal degree $j$ is $b^{1}_{j}$ (resp. $b^{0}_{j})$. We now compute:
-$b^{i}_{j}$ $=\dim_{k}\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)}\left(\Psi_{a}(\mathcal{C}),k(-j)[i-1]\right)$
-$=\dim_{k}\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)}\left(k(a-j),\Psi_{a}(\mathcal{C})[n-i]\right)$
-$=\dim_{k}\operatorname{Ext}^{n-i}_{X}\left(\Phi_{0}(k(a-j)),\mathcal{C}\right)$
-$=\dim_{k}\operatorname{Ext}_{X}^{r+a+1-i-2q}\left(i^{*}(\bigwedge^{r+a}\mathcal{T}_{\mathbb{P}^{n}_{k}})(-r-a),\mathcal{C}\right)$
-$=h^{r+a+1-i-2q}\left(X,\mathcal{C}\otimes i^{*}(\Omega^{r+a}_{\mathbb{P}^{n}_{k}})(r+a)\right)$
+Thus, the graded Betti number of $M$ in cohomological degree 0 (resp.
+1-1) and internal degree $j$ is $b^{1}_{j}$ (resp.
+$b^{0}_{j})$.
+By [Pav21, Lemma 3.1] and the identification of the shift $M[1]$ in $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ with the first syzygy of $M$, we have $b^{i}_{j}=\dim_{k}\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)}\left(\Psi_{a}(\mathcal{C}),k(-j)[i-1]\right)$.
+We now compute:
+[…]
+[…]
+[…]
+Finally, we have:
+$\operatorname{Ext}_{X}^{r+a+1-i-2q}\left(i^{*}(\bigwedge^{r+a}\mathcal{T}_{\mathbb{P}^{n}_{k}})(-r-a),\mathcal{C}\right)$ $=H^{r+a+1-i-2q}\left(X,\mathcal{C}\otimes i^{*}(\Omega^{r+a}_{\mathbb{P}^{n}_{k}})(r+a)\right)$
 $=H^{r+a+1-i-2q}\left(\mathbb{P}^{n}_{k},i_{*}(\mathcal{C})\otimes\Omega^{r+a}_{\mathbb{P}^{n}_{k}}(r+a)\right),$
-where the first equality follows from [Pav21, Lemma 3.1] and the identification of the shift $M[1]$ in $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ with the first syzygy of $M$, the second from Proposition 2.7, the third from Proposition 2.6, the fourth from Theorem 3.1, the fifth from a straightforward calculation, and the sixth from the projection formula.
+where the second equality follows from the projection formula.
 This proves (1).
-For (2), we have $\operatorname{rank}(F)=\operatorname{rank}(F^{0})+\operatorname{rank}(F^{1})=\sum_{j\in\mathbb{Z}}(b^{0}_{j}+b^{1}_{j})~{}=\rho(\mathcal{C}).$
+For (2), we note that $\operatorname{rank}(F)=\operatorname{rank}(F^{0})+\operatorname{rank}(F^{1})=\sum_{j\in\mathbb{Z}}(b^{0}_{j}+b^{1}_{j})~=\rho(\mathcal{C}).$
 Proof of Theorem 1.4.
@@
 Example 4.2.
-Let $\mathcal{F}$ be the structure sheaf of a point on $X$. We evidently have
-$\rho(\mathcal{F})=\sum_{r=0}^{n}\operatorname{rank}(\Omega^{r}_{\mathbb{P}^{n}_{k}})=2^{n+1}-1>2^{e+1}.$
+Let $\mathcal{F}$ be the structure sheaf of a point on $X$, and assume $n\geq 1$. We evidently have
+$\rho(\mathcal{F})=\sum_{r=0}^{n}\operatorname{rank}(\Omega^{r}_{\mathbb{P}^{n}_{k}})=2^{n}\geq 2^{e+1}.$
 Example 4.3.
@@
 For each $0\leq r\leq n$, there is a short exact sequence
-(4.5) $0\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r-d)\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to\mathcal{O}_{X}\otimes\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to 0$
+(4.5) $0\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r-d)\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to\mathcal{O}_{X}\otimes\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to 0.$
 Assume now that $a\leq 0$. By (4.4), we have:
@@
 Thus, each of the line bundles $\mathcal{O}_{X}(1-n),\dots,\mathcal{O}_{X}$ violates the inequality in Conjecture 1.3. Additionally, if $n=d=2$, then $\rho(\mathcal{O}_{X})=2<2^{e+1}=4$.
+As a concrete example, take $n=2$, j=1j=-1, and $d=1$. Formula 4.4 implies that
+$h^{i}(\mathbb{P}^{n}_{k},\Omega^{r}_{\mathbb{P}^{n}_{k}}(r-2))=\begin{cases}1,&r=i=2,\\ 0,&\text{else};\end{cases}\quad\text{and}\quad h^{i}(\mathbb{P}^{n}_{k},\Omega^{r}_{\mathbb{P}^{n}_{k}}(r-1))=\begin{cases}1,&r=i=1,\\ 0,&\text{else}.\end{cases}$
+It thus follows from the short exact sequence (4.5) (twisted by 1-1) that
+$\rho(\mathcal{O}_{X}(-1))=h^{1}(\mathbb{P}^{n}_{k},\mathcal{O}_{X}\otimes\Omega^{1}_{\mathbb{P}^{n}_{k}})+h^{1}(\mathbb{P}^{n}_{k},\mathcal{O}_{X}\otimes\Omega^{2}_{\mathbb{P}^{n}_{k}}(1))=h^{1}(\mathbb{P}^{n}_{k},\Omega^{1}_{\mathbb{P}^{n}_{k}})+h^{2}(\mathbb{P}^{n}_{k},\Omega^{2}_{\mathbb{P}^{n}_{k}})=2.$
+A nearly identical calculation shows that, when $n=d=2$, we have $\rho(\mathcal{O}_{X})=2$.
 Finally, we consider a family of examples where the bound in Conjecture 1.3 is achieved.
 Example 4.7.
-Assume $k$ contains a square root $i$ of $-1$, and suppose $f=x_{0}^{n}+\cdots+x_{n}^{n}$, where $n$ is even. In this case, $e=\frac{n}{2}$. We have graded matrix factorizations
+Assume $k$ contains a square root $i$ of 1-1, and suppose $f=x_{0}^{n}+\cdots+x_{n}^{n}$, where $n$ is even. In this case, $e=\frac{n}{2}$. We have graded matrix factorizations
 $F=\left(S(-e),S,x_{0}^{e},x_{0}^{e}\right)\in\operatorname{mf}_{\operatorname{gr}}(x_{0}^{n}),\quad F^{\prime}=\left(S(-e),S,x_{0}^{e}+ix_{1}^{e},x_{0}^{e}-ix_{1}^{e}\right)\in\operatorname{mf}_{\operatorname{gr}}(x_{0}^{n}+x_{1}^{n}).$